OpenStudy (anonymous):

Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary. A. 4/12 B. 7/4 C. 4/7 D. 12/4

hero (hero):

Hint: \[1 \frac{3}{4} = 1 + \frac{3}{4} = \frac{4}{4} + \frac{3}{4}\]

OpenStudy (anonymous):

7/4